# Planting and MCMC Sampling from the Potts model

### v1

@@ 2.3. Phases of the Potts model
Let $\nu=(\nu_{1},\ldots,\nu_{q})$ be a $q$-dimensional probability distribution. For an $n$-vertex graph $G$, let
@@
$Z_{\nu}(G)=\sum_{\sigma\in S_{\nu}}w_{G}(\sigma)$ and $\Psi=\Psi_{d,q,\beta}(\nu)=\lim_{\theta\downarrow 0}\lim_{n\rightarrow\infty}\tfrac{1}{n}\ln\mathbb{E}[Z_{\nu}(\boldsymbol{G})]$.
In short, the function $\Psi(\nu)$ captures the (logarithm of the) expected contribution to the partition function $Z(\boldsymbol{G})$ of the random graph $\boldsymbol{G}$ given by the configurations in $S_{\nu}$. As we shall see shortly, the local/global maximisers of $\Psi$ play a key role in determining the modes of the Gibbs distribution.
@@
which corresponds to the “disordered” configurations. As the temperature decreases and $\beta$ crosses the threshold $\beta_{u}$, another local maximum of $\Psi$ emerges at an “ordered” vector $\nu_{\mathrm{o}}$ where one of the $q$ colours dominates over the others (which are roughly balanced). More precisely, $\nu_{\mathrm{o}}$ is given by111 The value of $\beta_{u}$ is defined as the infinimum value of $\beta>0$ such that there exists $a\in(1/q,1)$ which satisfies (2.1); then it is simple to check that such an $a$ exists for all $\beta\geq\beta_{u}$. For certain values of $\beta$ (namely when $\beta\in(\beta_{u},\beta_{u}^{\prime})$), there may be more than one values of $a\in(1/q,1)$ that satisfy the equation; when this is the case, the value of $a$ that is relevant for the ordered phase is the largest such value.
@@
Interestingly, the local maximum at $\nu_{\mathrm{o}}$ does not become global until $\beta\geq\beta_{c}$ where $\beta_{c}=\ln\frac{q-2}{(q-1)^{1-2/d}-1}$ is known as the ordered/disordered threshold; notably, the only point where both $\nu_{\mathrm{o}}$ and $\nu_{\mathrm{d}}$ are global maxima is at $\beta=\beta_{c}$. Note also that $\nu_{\mathrm{d}}$ continues being a local maximum till $\beta<\beta_{u}^{\prime}$ where $\beta_{u}^{\prime}=\ln(1+\frac{q}{d-1})$; so, in the interval $(\beta_{u},\beta_{u}^{\prime})$ both $\nu_{\mathrm{d}}$ and $\nu_{\mathrm{o}}$ are local maxima of $\Psi$ which causes the metastability phenomena mentioned in the introduction, see [coja2023] for details.
For convenience, let $S_{\mathrm{d}}$ and $S_{\mathrm{o}}$ denote the sets of configurations $S_{\nu}$ when $\nu=\nu_{\mathrm{d}}$ and $\nu=\nu_{\mathrm{o}}$, respectively. For a graph $G$, define $Z_{\mathrm{d}}(G)$ and $Z_{\mathrm{o}}(G)$ analogously, and let $\mu_{G,\mathrm{d}}=\mu_{G}(\cdot\mid S_{\mathrm{d}})$ and $\mu_{G,\mathrm{o}}=\mu_{G}(\cdot\mid S_{\mathrm{o}})$ denoted the conditional Gibbs distributions. The following lemma from [BIShard] gives some sharp concentration estimates for $Z_{\mathrm{d}}(\boldsymbol{G})$ and $Z_{\mathrm{o}}(\boldsymbol{G})$ (derived using moments and the small subgraph conditioning method) and are the key behind our planting approach.
Lemma 2.1 ([BIShard, Theorem 5 & Lemma 16]).
Let $d,q\geq 3$ be integers, and $\beta\geq 0$ be real. Let $f(n)>0$ be any function with $f(n)=o(1)$. Whp over $\boldsymbol{G}$, we have:
[…]
Moreover, $\mu_{\boldsymbol{G}}\big{(}S_{\mathrm{d}}\big{)}=1-\mathrm{e}^{-\Omega(n)}$ when $\beta<\beta_{c}$, and $\mu_{\boldsymbol{G}}\big{(}\tilde{S}_{\mathrm{o}}\big{)}=1-\mathrm{e}^{-\Omega(n)}$ when $\beta>\beta_{c}$, where $\tilde{S}_{\mathrm{o}}$ is the union of $S_{\mathrm{o}}$ with its $q-1$ permutations. For $\beta=\beta_{c}$, it holds that $\mu_{\boldsymbol{G}}\big{(}S_{\mathrm{d}}\cup\tilde{S}_{\mathrm{o}}\big{)}=1-\mathrm{e}^{-\Omega(n)}$ and $\mu_{\boldsymbol{G}}(S_{\mathrm{d}}),\mu_{\boldsymbol{G}}(S_{\mathrm{o}})\geq f(n)$.
@@ 2.4. Planting disordered and ordered configurations
    $\beta<\beta_{u}^{\prime}$: if $\mathbb{P}\big{[}\big{(}\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}\big{)}\in{\mathcal{E}}\big{]}\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big{[}(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{d}}\big{)}\in{\mathcal{E}}\mid\boldsymbol{G}\big{]}\leq\frac{g(n)}{f(n)}$.
@@
    $\beta>\beta_{u}$: if $\mathbb{P}\big{[}\big{(}\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{o}}),\hat{\boldsymbol{\sigma}}_{\mathrm{o}}\big{)}\in{\mathcal{E}}\big{]}\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big{[}(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}}\big{)}\in{\mathcal{E}}\mid\boldsymbol{G}\big{]}\leq\frac{g(n)}{f(n)}$.
@@
Similarly, let $\rho^{G,\sigma}$ be the $[q]\times[q]$ matrix with the following half-edge colour statistics under $\sigma$, i.e.,
for colours $\mbox{for colours $i,j\in[q]$, $\rho^{G,\sigma}_{ij}=\frac{1}{2|E(G)|}\sum_{u,v\in V(G)}\mathbf{1}\big{\{}\{u,v\}\in E(G),\sigma(u)=i,\sigma(v)=j\big{\}}$}.$ (2.4)
@@
for $\mbox{for $i,j\in[q]$},\quad\rho_{\mathrm{d},ij}=\frac{\mathrm{e}^{\beta\mathbf{1}\{i=j\}}}{\sum_{i^{\prime},j^{\prime}\in[q]}\mathrm{e}^{\beta\mathbf{1}\{i^{\prime}=j^{\prime}\}}},\quad\rho_{\mathrm{o},ij}=\frac{\mathrm{e}^{\beta\mathbf{1}\{i=j\}}(\nu_{\mathrm{o},i}\nu_{\mathrm{o},j})^{(d-1)/d}}{\sum_{i^{\prime},j^{\prime}\in[q]}\mathrm{e}^{\beta\mathbf{1}\{i^{\prime}=j^{\prime}\}}(\nu_{\mathrm{o},i^{\prime}}\nu_{\mathrm{o},j^{\prime}})^{(d-1)/d}}.$ (2.5)
@@ 3.1. Phases for the random cluster model via Potts
For a graph $G\in\mathcal{G}_{n,d}$ and $\sigma\in[q]^{n}$, let $\boldsymbol{F}(G,\sigma)$ denote the random subset of edges obtained by keeping each monochromatic edge of $G$ under $\sigma$ with probability $p$. It is a standard fact (see, e.g., [RCMbook]) that, for integer $q\geq 3$, we can obtain a sample $\boldsymbol{F}\sim\pi_{G}$ by first sampling $\boldsymbol{\sigma}\sim\mu_{G}$ and then keeping each monochromatic edge of $G$ under $\sigma$ with probability $p=1-\mathrm{e}^{-\beta}$ (the so-called percolation step).
@@ 3.2. Mixing and Weak Spatial Mixing (WSM) within a phase
Treading a path initiated in [SinclairsGheissari2022], to prove Theorems 1.1 and Theorem 1.2 it suffices to show the following results for the disordered and ordered regimes, respectively.
@@
We first do the proof of Theorem 1.1. By Corollary 3.2, it holds that $\big{\|}\pi_{G}-\pi_{G,\mathrm{d}}\big{\|}_{\mathrm{TV}}=\mathrm{e}^{-\Omega(n)}$ and by Theorem 3.4 we can obtain a sample from $\pi_{G,\mathrm{d}}$ in $T$ steps using random-cluster dynamics. Then, using the translation between the Potts and RC models [RCMbook], this gives a sample $\hat{\boldsymbol{\sigma}}$ whose distribution is within TV-distance $\leq\varepsilon$ from $\mu_{G}$, as wanted. The proof of Theorem 1.2 is analogous using Corollary 3.2 and Theorem 3.5. ∎
It was shown in [SinclairsGheissari2022] that, for the ferromagnetic Ising model, a sufficient condition for fast convergence to equilibrium within the phase is the property of weak spatial mixing (WSM) within a phase, paired with some mixing-time estimates in local neighbourhoods. The same condition applies for the random-cluster model and, more generally, for monotone models. In the case of the random regular graph, the local neighbourhoods resemble the $d$-regular tree and the required mixing-time results have been established in [treeRC]. So, the main piece missing in order to obtain Theorems 3.4 and 3.5 is showing WSM within the corresponding phase. Roughly, the notion captures that, when we condition on the phase, the influence of a boundary configuration at distance $\ell$ is small.
More formally, for a graph $G=(V,E)$, an integer $\ell$ and a vertex $v$, let $B_{\ell}(v)$ be the graph induced by vertices at distance $\leq\ell$ from $v$, whose vertex and edge sets are going to be denoted by $V(B_{\ell}(v))$ and $E(B_{\ell}(v))$, respectively. A boundary condition $\tau$ on $B_{\ell}(v)$ is a subset of the edges in $E\backslash E(B_{\ell}(v))$; then $\pi^{\tau}_{G,v,\ell}$ denotes the RC distribution induced on $B_{\ell}(v)$ conditioned on $\tau$, i.e., for $S\subseteq E(B_{\ell}(v))$ it holds that $\pi^{\tau}_{G,v,\ell}(S)=\frac{\pi_{G}(S\cup\tau)}{\sum_{S^{\prime}\subseteq E(B_{\ell}(v))}\pi_{G}(S^{\prime}\cup\tau)}$. There are two extreme boundary conditions on $B_{\ell}(v)$ which will be most relevant: (i) the all-in or wired condition $+$ where all edges in $E\backslash E(B_{\ell}(v))$ are included in the boundary condition, and (ii) the all-out or free boundary condition -- where none of the edges in $E\backslash E(B_{\ell}(v))$ are included. We use $\pi^{+}_{G,v,\ell}$ and $\pi^{-}_{G,v,\ell}$ to denote the corresponding conditional RC distributions.
For an edge $e\in E$, let $1_{e}$ denote the subsets of $E$ that include $e$. Following [RCrandom], we say that we have WSM within the disordered phase (respectively, ordered) at distance $\ell$ if for each vertex $v$ and each edge $e$ incident to $v$ it holds that for each $\big{|}\pi^{-}_{G,v,\ell}(1_{e})-\pi_{G,\mathrm{d}}(1_{e})\big{|}\leq\frac{1}{100m}$ (respectively, $\big{|}\pi^{+}_{G,v,\ell}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big{|}\leq\frac{1}{100m}$) with $m=|E|$.
@@
Let $d\geq 3,q\geq(5d)^{5}$ be integers and $\beta\geq\beta_{c}$ be a real. Then, there exists $\delta>0$ such that, whp over $G\sim\mathcal{G}_{n,d}$, $G$ has WSM within the ordered phase at a distance $\ell=\lfloor(\frac{1}{2}-\delta)\log_{d-1}n\rfloor$. That is, with $m=|E|$, for every $v\in V$ and every edge $e$ incident to $v$, it holds that $\big{|}\pi^{+}_{G,v,\ell}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big{|}\leq\frac{1}{100m}$.
Using these, the proofs of Theorems 3.4 and 3.5 can be done using by now standard arguments; we give the details for completeness in Appendix E.

### v2

@@ 2.3. Phases of the Potts model
Let $\nu=(\nu_{1},\ldots,\nu_{q})$ be a $q$-dimensional probability vector. For an $n$-vertex graph $G$, let
@@
$Z_{\nu}(G)=\sum_{\sigma\in S_{\nu}}w_{G}(\sigma)$ and $\Psi=\Psi_{d,q,\beta}(\nu)=\lim_{\theta\downarrow 0}\lim_{n\rightarrow\infty}\tfrac{1}{n}\ln\mathbb{E}[Z_{\nu}(\boldsymbol{G})]$,
where $w_{G}(\sigma)=e^{\beta m(\sigma)}$ is the weight of $\sigma$. In short, the function $\Psi(\nu)$ captures the (logarithm of the) expected contribution to the partition function $Z(\boldsymbol{G})$ of the random graph $\boldsymbol{G}$ given by the configurations in $S_{\nu}$. As we shall see shortly, the local/global maximisers of $\Psi$ play a key role in determining the modes of the Gibbs distribution.
@@
which corresponds to the “disordered” configurations. As the temperature decreases and $\beta$ crosses the threshold $\beta_{u}$, another local maximum of $\Psi$ emerges at an “ordered” vector $\nu_{\mathrm{o}}$ where one of the $q$ colours dominates over the others (which are roughly balanced). More precisely, $\nu_{\mathrm{o}}$ is given by [^1] The value of $\beta_{u}$ is defined as the infinimum value of $\beta>0$ such that there exists $a\in(1/q,1)$ which satisfies (2.1); then it is simple to check that such an $a$ exists for all $\beta\geq\beta_{u}$. For certain values of $\beta$ (namely when $\beta\in(\beta_{u},\beta_{u}^{\prime})$), there may be more than one values of $a\in(1/q,1)$ that satisfy the equation; when this is the case, the value of $a$ that is relevant for the ordered phase is the largest such value.
@@
Interestingly, the local maximum at $\nu_{\mathrm{o}}$ does not become global until $\beta\geq\beta_{c}$ where $\beta_{c}=\ln\frac{q-2}{(q-1)^{1-2/d}-1}$ is known as the ordered/disordered threshold; notably, the only point where both $\nu_{\mathrm{o}}$ and $\nu_{\mathrm{d}}$ are global maxima is at $\beta=\beta_{c}$. Note also that $\nu_{\mathrm{d}}$ continues being a local maximum till $\beta<\beta_{u}^{\prime}$ where $\beta_{u}^{\prime}=\ln(1+\frac{q}{d-1})$; so, in the interval $(\beta_{u},\beta_{u}^{\prime})$ both $\nu_{\mathrm{d}}$ and $\nu_{\mathrm{o}}$ are local maxima of $\Psi$ which causes the metastability phenomena mentioned in the introduction, see [14] for details.
For convenience, let $S_{\mathrm{d}}$ and $S_{\mathrm{o}}$ denote the sets of configurations $S_{\nu}$ when $\nu=\nu_{\mathrm{d}}$ and $\nu=\nu_{\mathrm{o}}$, respectively. For a graph $G$, define $Z_{\mathrm{d}}(G)$ and $Z_{\mathrm{o}}(G)$ analogously, and let $\mu_{G,\mathrm{d}}=\mu_{G}(\cdot\mid S_{\mathrm{d}})$ and $\mu_{G,\mathrm{o}}=\mu_{G}(\cdot\mid S_{\mathrm{o}})$ denoted the conditional Gibbs distributions. The following lemma from [19] gives some sharp concentration estimates for $Z_{\mathrm{d}}(\boldsymbol{G})$ and $Z_{\mathrm{o}}(\boldsymbol{G})$ (derived using moments and the small subgraph conditioning method) and are the key behind our planting approach.
Lemma 2.1 ([19, Theorem 5 & Lemma 16]).
Let $d,q\geq 3$ be integers, and $\beta\geq 0$ be real. [^2] Here, we assume that the constant $\theta>0$, used in the definition of $Z_{\mathrm{d}}$ and $Z_{\mathrm{o}}$, is a sufficiently small constant depending only on $d,q,\beta$. Let $f(n)>0$ be any function with $f(n)=o(1)$. Whp over $\boldsymbol{G}$, we have:
 $\mbox{ $Z_{\mathrm{d}}(\boldsymbol{G})\geq f(n)\,\mathbb{E}[Z_{\mathrm{d}}(\boldsymbol{G})]$ if $\beta\leq\beta_{c}$}\quad\mbox{and}\quad\mbox{ $Z_{\mathrm{o}}(\boldsymbol{G})\geq f(n)\,\mathbb{E}[Z_{\mathrm{o}}(\boldsymbol{G})]$ if $\beta\geq\beta_{c}$}.$
Moreover, $\mu_{\boldsymbol{G}}\big(S_{\mathrm{d}}\big)=1-\mathrm{e}^{-\Omega(n)}$ when $\beta<\beta_{c}$, and $\mu_{\boldsymbol{G}}\big(\tilde{S}_{\mathrm{o}}\big)=1-\mathrm{e}^{-\Omega(n)}$ when $\beta>\beta_{c}$, where $\tilde{S}_{\mathrm{o}}$ is the union of $S_{\mathrm{o}}$ with its q1q-1 permutations. For $\beta=\beta_{c}$, it holds that $\mu_{\boldsymbol{G}}\big(S_{\mathrm{d}}\cup\tilde{S}_{\mathrm{o}}\big)=1-\mathrm{e}^{-\Omega(n)}$ and $\mu_{\boldsymbol{G}}(S_{\mathrm{d}}),\mu_{\boldsymbol{G}}(S_{\mathrm{o}})\geq f(n)$.
@@ 2.4. Planting disordered and ordered configurations
    $\beta\leq\beta_{c}$: if $\mathbb{P}\big[\big(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{d}}),\hat{\boldsymbol{\sigma}}_{\mathrm{d}}\big)\in{\mathcal{E}}\big]\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big[(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{d}}\big)\in{\mathcal{E}}\mid\boldsymbol{G}\big]\leq\frac{g(n)}{f(n)}$.
@@
    $\beta\geq\beta_{c}$: if $\mathbb{P}\big[\big(\hat{\boldsymbol{G}}(\hat{\boldsymbol{\sigma}}_{\mathrm{o}}),\hat{\boldsymbol{\sigma}}_{\mathrm{o}}\big)\in{\mathcal{E}}\big]\leq g(n)$, then, whp over $\boldsymbol{G}$, $\mathbb{P}\big[(\boldsymbol{G},\boldsymbol{\sigma}_{\boldsymbol{G},\mathrm{o}}\big)\in{\mathcal{E}}\mid\boldsymbol{G}\big]\leq\frac{g(n)}{f(n)}$.
@@
Similarly, let $\rho^{G,\sigma}$ be the $q]\times[q]$ matrix with the following half-edge colour statistics under $\sigma$, i.e.,
$\mbox{for colours $i,j\in[q]$, $\rho^{G,\sigma}_{ij}=\frac{1}{2|E(G)|}\sum_{u,v\in V(G)}\mathbf{1}\big\{\{u,v\}\in E(G),\sigma(u)=i,\sigma(v)=j\big\}$}.$ (2.4)
@@
$\mbox{for $i,j\in[q]$},\quad\rho_{\mathrm{d},ij}=\frac{\mathrm{e}^{\beta\mathbf{1}\{i=j\}}}{\sum_{i^{\prime},j^{\prime}\in[q]}\mathrm{e}^{\beta\mathbf{1}\{i^{\prime}=j^{\prime}\}}},\quad\rho_{\mathrm{o},ij}=\frac{\mathrm{e}^{\beta\mathbf{1}\{i=j\}}(\nu_{\mathrm{o},i}\nu_{\mathrm{o},j})^{(d-1)/d}}{\sum_{i^{\prime},j^{\prime}\in[q]}\mathrm{e}^{\beta\mathbf{1}\{i^{\prime}=j^{\prime}\}}(\nu_{\mathrm{o},i^{\prime}}\nu_{\mathrm{o},j^{\prime}})^{(d-1)/d}}.$ (2.5)
@@ 3.1. Phases for the random cluster model via Potts
For a graph $G\in\mathcal{G}_{n,d}$ and $\sigma\in[q]^{n}$, let $\boldsymbol{F}(G,\sigma)$ denote the random subset of edges obtained by keeping each monochromatic edge of $G$ under $\sigma$ with probability $p$. It is a standard fact (see, e.g., [24]) that, for integer $q\geq 3$, we can obtain a sample $\boldsymbol{F}\sim\pi_{G}$ by first sampling $\boldsymbol{\sigma}\sim\mu_{G}$ and then keeping each monochromatic edge of $G$ under $\sigma$ with probability $p=1-\mathrm{e}^{-\beta}$ (the so-called percolation step).
@@ 3.2. Mixing and Weak Spatial Mixing (WSM) within a phase
Treading a path initiated in [20], to prove Theorems 1.1 and Theorem 1.2 it suffices to show the following results for the disordered and ordered regimes, respectively.
@@
We first do the proof of Theorem 1.1. By Corollary 3.2, it holds that $\big\|\pi_{G}-\pi_{G,\mathrm{d}}\big\|_{\mathrm{TV}}=\mathrm{e}^{-\Omega(n)}$ and by Theorem 3.4 we can obtain a sample from $\pi_{G,\mathrm{d}}$ in $T$ steps using random-cluster dynamics. Then, using the translation between the Potts and RC models [24], this gives a sample $\hat{\boldsymbol{\sigma}}$ whose distribution is within TV-distance $\leq\varepsilon$ from $\mu_{G}$, as wanted. The proof of Theorem 1.2 is analogous using Corollary 3.2 and Theorem 3.5. ∎
It was shown in [20] that, for the ferromagnetic Ising model, a sufficient condition for fast convergence to equilibrium within the phase is the property of weak spatial mixing (WSM) within a phase, paired with some mixing-time estimates in local neighbourhoods. The same condition applies for the random-cluster model and, more generally, for monotone models. In the case of the random regular graph, the local neighbourhoods resemble the $d$-regular tree and the required mixing-time results have been established in [7]. So, the main piece missing in order to obtain Theorems 3.4 and 3.5 is showing WSM within the corresponding phase. Roughly, the notion captures that, when we condition on the phase, the influence of a boundary configuration at distance $\ell$ is small.
More formally, for a graph $G=(V,E)$, an integer $\ell$ and a vertex $v$, let $B_{\ell}(v)$ be the graph induced by vertices at distance $\leq\ell$ from $v$, whose vertex and edge sets are going to be denoted by $V(B_{\ell}(v))$ and $E(B_{\ell}(v))$, respectively. A boundary condition $\tau$ on $B_{\ell}(v)$ is a subset of the edges in $E\backslash E(B_{\ell}(v))$; then $\pi^{\tau}_{G,v,\ell}$ denotes the RC distribution induced on $B_{\ell}(v)$ conditioned on $\tau$, i.e., for $S\subseteq E(B_{\ell}(v))$ it holds that $\pi^{\tau}_{G,v,\ell}(S)=\frac{\pi_{G}(S\cup\tau)}{\sum_{S^{\prime}\subseteq E(B_{\ell}(v))}\pi_{G}(S^{\prime}\cup\tau)}$. There are two extreme boundary conditions on $B_{\ell}(v)$ which will be most relevant: (i) the all-in or wired condition ++ where all edges in $E\backslash E(B_{\ell}(v))$ are included in the boundary condition, and (ii) the all-out or free boundary condition - where none of the edges in $E\backslash E(B_{\ell}(v))$ are included. We use $\pi^{+}_{G,v,\ell}$ and $\pi^{-}_{G,v,\ell}$ to denote the corresponding conditional RC distributions.
For an edge $e\in E$, let $1_{e}$ denote the subsets of $E$ that include $e$. Following [18], we say that we have WSM within the disordered phase (respectively, ordered) at distance $\ell$ if for each vertex $v$ and each edge $e$ incident to $v$ it holds that $\big|\pi^{-}_{G,v,\ell}(1_{e})-\pi_{G,\mathrm{d}}(1_{e})\big|\leq\frac{1}{100m}$ (respectively, $\big|\pi^{+}_{G,v,\ell}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big|\leq\frac{1}{100m}$) with $m=|E|$.
@@
Let $d\geq 3,q\geq(5d)^{5}$ be integers and $\beta\geq\beta_{c}$ be a real. Then, there exists $\delta>0$ such that, whp over $\sim\mathcal{G}_{n,d}, GG$ has WSM within the ordered phase at a distance $\ell=\lfloor(\frac{1}{2}-\delta)\log_{d-1}n\rfloor$. That is, with $m=|E|$, for every $v\in V$ and every edge $e$ incident to $v$, it holds that $\big|\pi^{+}_{G,v,\ell}(1_{e})-\pi_{G,\mathrm{o}}(1_{e})\big|\leq\frac{1}{100m}$.
Using these, the proofs of Theorems 3.4 and 3.5 can be done using by now standard arguments; we give the details for completeness in Appendix F.